# Information-theoretic description of a feedback-control Kuramoto model

### v1

@@ III.1.2 Agent Description
The agent is a localized, mobile perturbation whose purpose is to maintain a global desynchronized state. The agent’s degrees of freedom are $a=(i,\kappa_{i})$, where $i$ is the index of the node it is at, and $\kappa_{i}\in\{\pm 1\}$ is the direction of a perturbation of strength $\alpha$ it applies to the oscillator at $i$. After perturbing, the agent jumps to an adjacent node uniformly at random; the entire jump-perturbation action takes a time much shorter than the environmental timescale, $\tau\ll 1$, but not so fast that it can perturb more than a small fraction of the network on that same timescale, $\tau\gg 1/N$. This gives a natural way to parametrize the speed of the agent as $\beta=1/N\tau\ll 1$. As we will show, the single parameter $\alpha\beta$, fully describes the behavior of an agent. A schematic of the process is shown in Fig. 1.
@@ III.1.3 Dynamics
where $k_{i}=\sum_{j}A_{ij}$ is the degree of node $i$. In the fully-connected Kuramoto model, $\sigma=K/N$, and Eq.(1) describes a set of oscillators that interact solely through their mutual coupling to the mean field,
$\frac{d\theta_{i}}{dt}=\omega_{i}-Kr\sin(\theta_{i}-\phi).$ (4)
For arbitrary network topologies one can show that Eq.(1) can be written in an analogous way using the local order parameters Rodrigues et al. [2016],
@@
In the thermodynamic limit, $N\rightarrow\infty$, in the fully-connected model, above a critical coupling ($K_{c}=2/\pi g(0)$) oscillators begin to synchronize, $r>0$. Analysis of mean field formulations for more complex topologies suggest that $\sigma_{c}=K_{c}\langle k\rangle/\langle k^{2}\rangle$, where $\langle k^{n}\rangle$ is the $n^{\text{th}}$ moment of the degree distribution of $G$. Numerical results suggest a more complicated relationship with topology, while confirming that $\sigma_{c}\propto K_{c}$. In our model the coupling is driven exogenously, that is perturbations to the system from outside of itself can drive $\sigma$ to values above and below $\sigma_{c}$ (for instance exposure to stroboscopic lights triggering an epileptic fit, or magnetic fields altering cardiac rhythm). Of interest to us will be when the system transitions through the critical coupling from below, shifting the system equilibrium to synchronized states. As the environment slides towards its new equilibrium, the agent as a control process, seeks to maintain a desynchronized state. Thus it must tune its actions towards that goal, while only using local information, that is knowledge only of the location $i$ (the node) and its immediate neighborhood.
@@ III.1.4 Agent-effect on synchronization dynamics
We simulate an ensemble of agents with $\alpha=0.5$ and $\beta=0.01$ on two types of graphs, with the same average degree $\langle k\rangle$ but increasing degree-variance $\sigma_{k}^{2}=\langle k^{2}\rangle-\langle k\rangle^{2}$. (See Appendix A for details of the simulation.) In Figure  2 we plot the global order parameter $\mathds{E}[r]$, as a function of the coupling strength $\sigma$. The dashed curves represent the Kuramoto environment without the agent, whereas the solid curves represent agent incorporation. In all three cases we see the effect of the agent is to delay the onset of synchronization, with the most dramatic suppression seen in the scale-free graph. Simulations varying $\alpha,\beta$ exhibit strikingly similar behavior between an $(\alpha,\beta)$ agent and a $(\alpha^{\prime},\beta^{\prime})$ agent if $\alpha\beta=\alpha^{\prime}\beta^{\prime}$.
Figure 2: The expected global order parameter for Poisson (top), and scale-free (bottom) graphs ($N=10^{3}$) as a function of the coupling strength $\sigma$. All networks have an average degree $\langle k\rangle=4$. The dashed curves represent a pure Kuramoto environment, while the solid curves represent the incorporation of the intervening agent. In both cases the shaded regions represents 1-$\sigma$ fluctuations from an ensemble computation. Note that the vertical axis scales logarithmically on all the plots, while the horizontal axis is linear for the top plot and logarithmic on the bottom one.
@@
where $W_{i}$ is a Wiener process. Incorporation of an agent leads to the addition of two dynamical terms in Eq. (1) parametrized by the perturbation strength and timescale $\alpha$ and $\beta$, respectively. Both parameters enter in solely through their product, providing an explanation for the observed degeneracy.
The two additional terms in Eq. (III.1.4) have interesting physical interpretations. The second one plays the role of a heat bath as in other variants of the Stochastic Kuramoto model Sonnenschein et al. [2013]; Sonnenschein and Schimansky-Geier [2013]. The key difference here is that the strength of the fluctuations at location $i$ is modulated by the degree of the node $k_{i}$. Instead of a single heat bath, there are several at temperatures quantized in units of $1/\langle k\rangle$. Nodes with higher degree are coupled to hotter baths, while those with lesser degree to colder ones.
@@
$\omega_{i}\mapsto\begin{cases}\omega_{i}+\alpha\beta\frac{k_{i}}{\langle k\rangle}&\omega_{i}>0,\\ \omega_{i}-\alpha\beta\frac{k_{i}}{\langle k\rangle}&\omega_{i}<0.\\ \end{cases}$ (8)
This shift changes the synchronization criterion to $|\omega_{i}|<\sigma k_{i}r_{i}-\alpha\beta k_{i}\langle k\rangle$. Oscillators with positive natural frequencies between $\sigma k_{i}r_{i}-\alpha\beta k_{i}/\langle k\rangle<\omega_{i}<\sigma k_{i}r_{i}$ or negative natural frequencies between $-\sigma k_{i}r_{i}<\omega_{i}<-\sigma k_{i}r_{i}+\alpha\beta k_{i}/\langle k\rangle$ that were synchronized in the absence of an agent will desynchronize in the presence of one. Consequently the local order parameters, $r_{i}$, decreases, causing a cascade of oscillators inside the old synchronization range to follow suit until a new equilibrium is reached.
The combined effect of the two agent terms in Eq.  (III.1.4) can best be understood qualitatively via their effect on the frequency distribution, $g(\omega)$. The heat bath alone acts as a Gaussian convolution kernel, which is then shifted as per Eq. (8). For any reasonable initial distribution $g$ convolved with the kernel $h$, results in a deformed distribution, $g^{A}=h*g$ of higher variance (* is the convolution operator). This deformation stretches the distribution and, to conserve probability, compresses its height such that $g^{A}(0)<g(0)$. Given that $\sigma_{c}^{A}/\sigma_{c}=K_{c}^{A}/K_{c}=g(0)/g^{A}(0)>1$, the critical coupling in the presence of the agent increases which, in turn, delays the onset of synchronization.
@@
The full state space of our system is $S=(-\pi,\pi]^{N}\times\{1,2,\dots,N\}\times\{\pm 1\}$. The last of these describe the agent’s action $\kappa_{i}$, based on their observation of the local mean field $\hat{\phi}_{i}$. Combining this with $\theta_{i}$, the agent determines whether oscillator $i$ is running ahead or behind the mean field of its neighbors. For the agent, the state space of the environment is coarse grained to a string of bits ${\boldsymbol{b}}$. The action of the agent depends on the outcome of its measurement, and the maximal mutual information between the environment and the agent’s action degree of freedom is one bit per jump. The $i^{\text{th}}$ bit $b_{i}$ is $1$ if the oscillator at the $i^{th}$ node is running ahead of its mean field, or 00 otherwise.
@@
As $\sigma$ approaches the critical coupling $\sigma_{c}$, there is a slowing down of the computation due to the divergence of the time-scale $\tau\sim(\sigma_{c}-\sigma)^{-1/2}$. As the equation indicates, nodes with high connectivity (i.e hubs) have the longest bit-flip timescales, whereas lower-degree nodes compute on a faster time-scale.
@@ III.3 Viability Measure $\mathcal{V}$
Figure 3: The efficacy, Eq. (10) (solid black lines), for agents on Poisson graphs (top panel) and scale-free graphs (bottom panel) with $\alpha=0.5$ and $\beta=0.01$. The noised efficacies, Eq. (23), are shown as colored curves with blue corresponding to low noise and pink to high noise. The curves are averaged over an ensemble of 128 graphs for each value of coupling and noise.
The efficacy is the relative difference between the expected value of the synchronicity of the environment including the agent ($\mathds{E}[r|\alpha,\beta]$) and without it ($\mathds{E}[r]$) thus,
@@
In Fig. (3), we plot the efficacy (Eq. (10) as solid black lines in function of $\sigma$ for both types of networks. In all cases we see a non-monotonic behavior, with a peak in $\mathcal{E}$ at intermediate values of $\sigma$. Qualitatively speaking, for low values of the coupling ($\sigma<<\sigma_{c}$) the effect of the agent is minimal given that the system is in an asychronous state and therefore there is no need to intervene. On the other hand in the strong coupling regime $\sigma>>\sigma_{c}$, the agent also has low efficacy—given that $\beta<<1$, a large fraction of nodes synchronize at a timescale faster than its ability to influence these nodes. Thus there is a peak in viability at intermediate values of $\sigma$ where the agent can exert its strongest influence in preventing the synchronization of a non-negligible fraction of oscillators.
It is instructive to check whether the peak in efficacy occurs at (or near) the critical coupling $\sigma_{c}$. While an analytical treatment for complex topologies is difficult, progress can be made by resorting to the fully-connected version of the Kuramoto model, for which $\sigma=K$. As noted in Sec. III.1 the $\alpha\beta$-agent stretches the frequency distribution $g(\omega)$, increasing its variance from unity to $1+\alpha\beta$. The derivatives of the frequency distribution at $\omega=0$, namely $g^{(2n)}(0)$ (see appendix C for details), define the critical coupling and the behavior of the order parameter near it,
$r(K)=2\sqrt{2\frac{g(0)}{|g^{\prime\prime}(0)|}}\frac{(K-K_{c})^{1/2}}{K^{3/2}}.$ (12)
We next calculate the effect of the change in variance of $g(\omega)$ on these derivatives.
Denoting the stretched distribution $\tilde{g}(\omega)$, a Gaussian with variance $1+\alpha\beta$; we relate it to the original distribution in the limit $\alpha\beta\ll 1$:
$\tilde{g}(\omega)$ $=\frac{e^{-\frac{1}{2(1+\alpha\beta)}\omega^{2}}}{\sqrt{2\pi(1+\alpha\beta)}}$
$\approx\left(1+\frac{\omega^{2}-1}{2}\alpha\beta\right)g(\omega).$ (13)
The critical coupling of the deformed distribution, $\tilde{K}_{c}$, is consequently shifted according to,
$\tilde{K}_{c}=\frac{2}{\pi\tilde{g}(0)}\approx K_{c}\left(1+\frac{\alpha\beta}{2}\right).$ (14)
@@
For environments with coupling below $K_{c}$, the agent is not effective because the environment dynamics are already driven towards asynchronous states. However, for $K_{c}<K<\tilde{K}_{c}$, the efficacy jumps to $1$ since $\mathds{E}_{\alpha\beta}[r]$ vanishes while $\mathds{E}[r]$ does not. At larger couplings the agent+environment synchronization begins to drop. To first order in $\alpha\beta$, the efficacy can be approximated as
@@
$\mathcal{E}\approx\left\{\begin{array}[]{cc}0&K<K_{c}\\ 1&K_{c}\leq K<\tilde{K}_{c}\\ \frac{\alpha\beta}{4}\frac{1}{\frac{K}{K_{c}}-1}&K\geq\tilde{K}_{c}.\end{array}\right.$ (18)
@@
Thus, as also seen in the simulations of more complex topologies in Fig. 3, the efficacy is peaked above the critical coupling of the agent-free environment $\sigma_{c}$, and decreases beyond the critical coupling of the agent-inclusive environment $\tilde{\sigma}_{c}$. Based on these results we are now able to both define viability and characterize how it responds to agent dynamics. To derive semantic information content in the system we now explore the architecture of correlations in the agent/environment coupling.
@@
With the computational description in mind, a natural way of scrambling presents itself. A noisy channel reduces the information transmission rate, and therefore, the control process will not be able to modulate its behavior in a way that reduces synchronization in the background process, thus affecting its viability. Consequently, scrambling the correlations between agent and environment can be implemented by introducing noise in the agent’s information about the phase of the neighborhood of node $i$ that it visits.
Let this noise be parameterized by $\eta\in[0,1]$, where $\eta=0$ corresponds to the agent with $\hat{\phi}_{i}=\phi_{i}$, and $\eta=1$ an agent for whom $\hat{\phi}_{i}$ is drawn uniform randomly between $-\pi$ and $\pi$. For values $10<\eta<$ 1, the estimator is a random variable $\hat{\phi}_{i}\sim\mathcal{U}(\phi_{i}-\eta\pi,\phi_{i}+\eta\pi)$. Then for node $i$, defining $\eta_{c}=\min(|\theta_{i}-\phi_{i}|/\pi,1-|\theta_{i}-\phi_{i}|/\pi)$, the probability that the agent’s kick is in the wrong direction is
$p(\hat{\kappa}_{i}\neq\kappa_{i}|\phi_{i},\theta_{i})=\left\{\begin{array}[]{ll}0&\text{if }\eta\leq\eta_{c}\\ \frac{1}{2}(1-\frac{\eta_{c}}{\eta})&\text{if }\eta_{c}<\eta<1-\eta_{c}\\ 1-\frac{1}{2\eta}&\text{otherwise}\end{array}\right..$ (22)
@@
Averaged over all nodes visited by the agent per unit time, the contribution due to noising from both synced and unsynced oscillators further reduces the information transmission rate from environment to agent. To quantify the effect of scrambling on the viability we modify Eq. (10) by defining the noised efficacy
@@
We plot the results of introducing noise in Fig. (3) which shows the change in efficacy (colored solid lines) in the range $0\leq\eta\leq 1$. For both networks, the peak efficacy is reduced by an order of magnitude for maximal noising; while the overall shape of the curve is preserved with increased noise the curve is increasingly flattened. Note that even at maximal noising, where the agent has no correlations with the environment, the efficacy is still finite. This is easy to understand in terms of the agent acting as a heat bath—even though it is no longer making an informed choice in the way it kicks its local oscillator, the net effect is still to deform the frequency distribution enough to increase the critical coupling to $\tilde{\sigma}_{c}>\sigma_{c}$, and thus delay the onset of synchronization.
Figure 4: The noised efficacy Eq. (23) (top panels) as a function of unscrambled information (measured in bits) and coupling $\sigma$ for Poisson graphs (left), and scale-free graphs (right). The noised viability Eq. (24) is plotted in the bottom panels as a function of unscrambled information. The pink dashed lines in these are the best fit linear models, giving a viability per bit (VpB) of $0.171\pm 0.002$, and $0.209\pm 0.003$ for Poisson and scale-free graphs respectively.
@@
Without loss of generality, we assume a uniform $p(\sigma)$ with support $\sigma\in[0,1]$ when computing the viabilities. The efficacy as a function of both scrambling (measured as amount of unscrambled bits) and the coupling is shown in the top panels of Fig. 4, with the light blue regions corresponding to those of high efficacy. The width of the regions is highest for low scrambling, and it tapers as the amount of scrambled information increases. We note that in both cases, the region of peak efficacy is relatively unchanged with increasing noise $\eta$. The high-efficacy region occurs for lower coupling strength in the scale-free case, but this is to be expected since $\sigma_{c}^{\textrm{scale-free}}<\sigma_{c}^{\textrm{Poisson}}$ as $\langle k^{2}\rangle^{\textrm{scale-free}}>\langle k^{2}\rangle^{\textrm{Poisson}}$.

### v3

@@ III.1.2 Agent Description
The agent is a localized, mobile perturbation whose purpose is to maintain a global desynchronized state. The agent’s degrees of freedom are $a=(i,\kappa_{i})$, where $i$ is the index of the node it is at, and $\kappa_{i}\in\{\pm 1\}$ is the direction of a perturbation of strength $\alpha$ it applies to the oscillator at $i$. After perturbing, the agent jumps to an adjacent node uniformly at random; the entire jump-perturbation action takes a time much shorter than the oscillator timescale, $\tau_{J}\ll 1$, but not so fast that it can perturb more than a small fraction of the network on that same timescale, $\tau_{J}\gg 1/N$. This gives a natural way to parameterize the speed of the agent as $\beta=(N\tau_{J})^{-1}\ll 1$. As we will show, the single parameter $\alpha\beta$, fully describes the behavior of an agent. A schematic of the process is shown in Fig. 1.
Figure 1: Oscillators $\{\theta_{i}\}$ with natural frequencies $\{\omega_{i}\}$ are placed on the nodes of different types of networks; shown here power law (upper left) and Poisson graphs (upper right). Kuramoto dynamics are imposed resulting in Kuramoto environments. At each node (lower left) there is a local order parameter and local mean field, $\{r_{i},\phi_{i}\}$. Together, these dynamical quantities define the local mean field coupling for each oscillator. Control agents are placed in the Kuramoto environments, moving uniformly at random along edges, and perturbing the nodes based off whether the local mean field estimate is running ahead or behind the oscillator, $\hat{\kappa}_{i}=\text{sign}(\phi_{i}-\theta_{i})$. Here $\alpha$ describes the strength of the perturbation that the agent applies, and $\beta$ the time-scale or speed of the perturbation.
@@ III.1.3 Dynamics
where $k_{i}=\sum_{j}A_{ij}$ is the degree of node $i$. In the fully-connected Kuramoto model Eq.(1) describes a set of oscillators that interact solely through their mutual coupling to the mean field,
$\frac{d\theta_{i}}{dt}=\omega_{i}-Kr\sin(\theta_{i}-\phi),$ (4)
with $K$ described earlier. For arbitrary network topologies Eq.(1) can be written in an analogous way using the local order parameters Rodrigues et al. [2016],
@@
In the thermodynamic limit, $N\rightarrow\infty$, of the fully-connected model, above a critical coupling ($K_{c}=2/\pi g(0)$) oscillators begin to synchronize, $r>0$. Analysis of mean field formulations for more complex topologies suggest that $\sigma_{c}=K_{c}\langle k\rangle/\langle k^{2}\rangle$, where $\langle k^{n}\rangle$ is the $n^{\text{th}}$ moment of the degree distribution of $G$. Numerical results suggest a more complicated relationship with topology, while confirming that $\sigma_{c}\propto K_{c}$. In our model the coupling is driven exogenously, that is perturbations to the system from outside of itself can drive $\sigma$ to values above and below $\sigma_{c}$ (for instance exposure to stroboscopic lights triggering an epileptic fit, or magnetic fields altering cardiac rhythm). Of interest to us will be when the system transitions through the critical coupling from below, shifting the system equilibrium to synchronized states. As the environment slides towards its new equilibrium, the agent as a control process, seeks to maintain a desynchronized state. Thus it must tune its actions towards that goal, while only using local information, that is knowledge gained only at the location $i$ (the node) and its immediate neighborhood.
@@ III.1.4 Agent-effect on synchronization dynamics
We simulate an ensemble of agents with $\alpha=0.5$ and $\beta=0.16$, first on the fully-connected version of the model, and then on two types of graphs, with the same average degree $\langle k\rangle$ but increasing degree-variance $\langle k^{2}\rangle-\langle k\rangle^{2}$. (See Appendix A for details of the simulation.) In Figure  2 we plot the global order parameter $\mathds{E}[r]$, as a function of the coupling strength $\sigma$. The dashed curves represent the Kuramoto environment without the agent, whereas the solid curves represent agent incorporation. In all cases, we see the effect of the agent is to delay the onset of synchronization, with the most dramatic suppression seen in fully-connected and scale-free graphs. Simulations varying $\alpha,\beta$ exhibit identical behavior between an $(\alpha,\beta)$ agent and a $(\alpha^{\prime},\beta^{\prime})$ agent if $\alpha\beta=\alpha^{\prime}\beta^{\prime}$.
Figure 2: The expected global order parameter for fully-connected (top), Poisson (middle), and scale-free (bottom) graphs ($N=10^{3}$) as a function of the coupling strength. For the fully connected version, shown as a vertical dashed blue line is the critical coupling $K_{c}$ in the absence of an agent, and in dashed vertical pink line, the the approximation of the critical coupling in the presence of an agent, $K^{A}_{c}$. All other networks have an average degree $\langle k\rangle=4$. The dashed black curves represent a pure Kuramoto environment, while the solid black curves represent incorporation of the intervening agent. In both cases the shaded regions represents one standard deviation fluctuations from an ensemble computation. Note that the vertical axis scales linearly in the top plot, and logarithmically in the bottom two plots, while the horizontal axis is linear for the top two plots and logarithmic on the bottom one.
@@
where $W_{i}$ is a Wiener process. Incorporation of an agent leads to the addition of two dynamical terms in Eq. (1) parameterized by the perturbation strength and timescale $\alpha$ and $\beta$, respectively. Both parameters enter in solely through their product, providing an explanation for the observed degeneracy.
The two additional terms in Eq. (III.1.4) have interesting physical interpretations. The second one plays the role of a heat bath as in other variants of the Stochastic Kuramoto model Sonnenschein et al. [2013]; Sonnenschein and Schimansky-Geier [2013]; Sakaguchi and Kuramoto [1986]; Sakaguchi et al. [1988]. The key difference here is that the strength of the fluctuations at location $i$ is modulated by the degree of the node $k_{i}$. Instead of a single heat bath, there are several at temperatures quantized in units of $1/\langle k\rangle$. Nodes with higher degree are coupled to hotter baths, while those with lesser degree to colder ones.
@@
$\omega_{i}\mapsto\tilde{\omega}_{i}=\begin{cases}\omega_{i}+\alpha\beta\frac{k_{i}}{\langle k\rangle}&\omega_{i}>0,\\ \omega_{i}-\alpha\beta\frac{k_{i}}{\langle k\rangle}&\omega_{i}<0.\\ \end{cases}$ (8)
This shift changes the synchronization criterion to $|\omega_{i}|<\sigma k_{i}r_{i}-\alpha\beta k_{i}/\langle k\rangle$. Oscillators with positive natural frequencies between $\sigma k_{i}r_{i}-\alpha\beta k_{i}/\langle k\rangle<\omega_{i}<\sigma k_{i}r_{i}$ or negative natural frequencies between $-\sigma k_{i}r_{i}<\omega_{i}<-\sigma k_{i}r_{i}+\alpha\beta k_{i}/\langle k\rangle$ that were synchronized in the absence of an agent will desynchronize in the presence of one. Consequently the local order parameters, $r_{i}$, decreases, causing a cascade of oscillators inside the old synchronization range to follow suit until a new equilibrium is reached.
The combined effect of the two agent terms in Eq.  (III.1.4) can best be understood qualitatively via their effect on the frequency distribution, $g(\omega)$. The shift produces a new distribution $\tilde{g}(\tilde{\omega})=g(\omega)$, which vanishes for $|\tilde{\omega}_{i}|<\alpha\beta k_{i}/\langle k\rangle$ In the thermodynamic limit, for a random node in the network $\omega$ is a random variable drawn from this distribution. The heat bath introduces a fluctuation to the frequency, another random variable, $\delta\omega$. The distribution of $\omega+\delta\omega$, the sum of two random variables, is the convolution of their distributions, so the end result is the effective distribution $g^{A}(\omega)=(\tilde{g}*h_{\alpha\beta})(\omega)$, where $h_{\alpha\beta}(\delta\omega)$ is a Gaussian with vanishing mean and variance $\alpha\beta$ Widder and Hirschman [2015]. Both operations increase the variance, so $g^{A}$ is a stretched version of $g$ with a suppressed peak. Since the critical point depends inversely on this value, this will necessarily increase the critical coupling.
@@
The full state space of our system is $S=(-\pi,\pi]^{N}\times\{1,2,\dots,N\}\times\{\pm 1\}$. The last of these describe the agent’s action $\kappa_{i}$, based on their observation of the local mean field $\hat{\phi}_{i}$. Combining this with $\theta_{i}$, the agent determines whether oscillator $i$ is running ahead or behind the mean field of its neighbors. For the agent, the state space of the environment is coarse grained to a string of bits ${\boldsymbol{b}}$. The $i^{\text{th}}$ bit $b_{i}$ is $1$ if the oscillator at the $i^{th}$ node is running ahead of its mean field, or 00 otherwise. The action of the agent depends on the outcome of its measurement, and the maximal mutual information between the environment and the agent’s action degree of freedom is one bit per jump.
@@
As $\sigma$ approaches the critical coupling $\sigma_{c}$, there is a slowing down of the computation due to the divergence of the time-scale $\tau_{f}\sim(\sigma_{c}-\sigma)^{-1/2}$. As the equation indicates, nodes with high connectivity (i.e hubs) have the longest bit-flip timescales, whereas lower-degree nodes compute on a faster time-scale.
@@ III.3 Viability Measure $\mathcal{V}$
Figure 3: The efficacy, Eq. (10) (solid black lines), for agents on fully-connected (top), Poisson (middle), and scale-free graphs (bottom) graphs with $\alpha=0.5$ and $\beta=0.16$. The fully connected curves are averaged over ensembles of $1024$ different graphs; the other two have ensemble sizes of $128$. The noised efficacies, Eq. (20), are shown as colored curves with blue corresponding to low noise and pink to high noise. For fully-connected graphs the ensemble is of size $1024$ again, while the Poisson and scale-free graphs have smaller ensembles of $32$. The dashed black line in the top panel is our analytical approximation  15 of the agent efficacy.
The efficacy is the relative difference between the expected value of the synchronicity of the environment including the agent, $\mathds{E}[r|\alpha,\beta]$, and without it, $\mathds{E}[r]$, thus,
@@
In Fig. (3), we plot the efficacy (Eq. (10) as solid black lines in function of $\sigma$ for the three types of networks. In all cases we see a non-monotonic behavior, with a peak in $\mathcal{E}$ at intermediate values of $\sigma$. Qualitatively speaking, for low values of the coupling ($\sigma\ll\sigma_{c}$) the effect of the agent is minimal given that the system is in an asynchronous state and therefore there is no need to act. On the other hand in the strong coupling regime $\sigma\gg\sigma_{c}$, the agent also has low efficacy—given that $\beta\ll 1$, a large fraction of nodes synchronize at a timescale faster than its ability to influence these nodes. Thus there is a peak in viability at intermediate values of $\sigma$ where the agent can exert its strongest influence in preventing the synchronization of a non-negligible fraction of oscillators.
It is instructive to check the fully-connected model, where $\sigma=K/N$, to see if we can make further progress. In the thermodynamic limit, an exact calculation for the critical coupling for the agent-free environment leads to $K_{c}=1.595$, plotted as a blue dashed line in the upper panel of Fig. 2. For the agent-based environment, through a sequence of manipulations, one can arrive at a numerical estimate for the shifted critical coupling $K^{A}_{c}\approx 1.988$, plotted as a pink dashed line in the same figure. (See Appendices  C and  D for details of the calculation, including analytically tractable estimates for $K^{A}_{c}$.)
@@
To estimate the efficacy, we note that in environments with coupling below $K_{c}$, we expect the agent to be ineffective given the environmental dynamics are driven towards asynchronous states. Above $K_{c}$ but below $K^{A}_{c}$, the efficacy should jump to $1$ since one expects $\mathds{E}_{\alpha\beta}[r]$ to vanish while $\mathds{E}[r]$ does not. At couplings larger than $K_{c}^{A}$ the agent+environment synchronization. To first order in $\alpha\beta$, the efficacy can be approximated as
@@
$\mathcal{E}\approx\left\{\begin{array}[]{cc}0&K<K_{c}\\ 1&K_{c}\leq K<K^{A}_{c}\\ \frac{\alpha\beta}{4}\frac{K_{c}}{K-K_{c}}&K>K^{A}_{c}.\end{array}\right.$ (15)
This is plotted in the top panel of figure 3, where it captures many of the features of the actual efficacy. We note that our estimate differs in the $K<K_{c}$ regime. This is unexpected because as $K\rightarrow 0$ the distribution of phases should approach a Rayleigh distribution describing the magnitude of the sum of a set of randomly directed unit vectors. In hindsight, this is not hard to understand. The Rayleigh distribution takes into account even those configurations where the random vectors all point in similar directions. The agent disrupts these types of arrangements, shifting probability from the tail of the Rayleigh distribution, thereby lowering the expected value of the magnitude i.e. the global order parameter.
@@
With the the fully connected model as reference, similar features are found in the simulations of more complex topologies in Fig. 3; the efficacy is peaked above the critical coupling of the agent-free environment $\sigma_{c}$, and decreases beyond the critical coupling of the agent-inclusive environment $\sigma^{A}_{c}$. Based on these results we are now able to both define viability and characterize how it responds to agent dynamics. To derive semantic information content in the system we now explore the architecture of correlations in the agent/environment coupling.
@@
With the computational description in mind, a natural way of scrambling presents itself. A noisy channel reduces the information transmission rate, and therefore, the control process will not be able to modulate its behavior in a way that reduces synchronization in the background process, thus affecting its viability. Consequently, scrambling the correlations between agent and environment can be implemented by introducing noise in the agent’s information about the phase of the neighborhood of node $i$ that it visits. We interpret this scrambling as a decrease in the fidelity of the sensor used by the agent to probe the mean field.
@@
$p(\hat{\kappa}_{i}\neq\kappa_{i}|\phi_{i},\theta_{i})=\left\{\begin{array}[]{ll}0&\text{if }\eta\leq\eta_{c}\\ \frac{1}{2}(1-\frac{\eta_{c}}{\eta})&\text{if }\eta_{c}<\eta<1-\eta_{c}\\ 1-\frac{1}{2\eta}&\text{otherwise.}\end{array}\right.$ (19)
[…]
Averaged over all nodes visited by the agent per unit time, the contribution due to noising from both synchronized and unsynchronized oscillators further reduces the information transmission rate from environment to agent. To quantify the effect of scrambling on the viability we modify Eq. (10) by defining the noised efficacy
@@
We plot the results of introducing noise in Fig. (3) which shows the change in efficacy (colored solid lines) in the range $0\leq\eta\leq 1$. For both networks, the peak efficacy is reduced by an order of magnitude for maximal noising; while the overall shape of the curve is preserved with increased noise the curve is increasingly flattened. Note that even at maximal noising, where the agent has no correlations with the environment, the efficacy is still non-zero. This is easy to understand in terms of the agent acting as a heat bath—even though it is no longer making an informed choice in the way it kicks its local oscillator, the net effect is still to deform the frequency distribution enough to increase the critical coupling to $\sigma^{A}_{c}>\sigma_{c}$, and thus delay the onset of synchronization.
Figure 4: The noised efficacy Eq. (20) (top panels) as a function of sensor fidelity (horizontal axis) and coupling strength,$K$, (vertical axis) for fully-connected graphs (left) and $\sigma$ for Poisson graphs (middle), and scale-free graphs (right). For the fully-connected the horizontal dashed pink line is the analytical value of $K_{c}$. In all three panels the dashed black line traces the maximal efficacy at each value of sensor fidelity. The noised viability Eq. (21), scaled to the actual viability is plotted in the lower two rows of panels as a function of unscrambled information. The middle row assumes coupling strengths are drawn from a uniform distribution, while the bottom row has them drawn from three different exponential distributions. The pink dashed lines in these are the best fit linear models, with further details in the text.
@@
In principle, there is no restriction on the distribution $\rho(\sigma)$. For illustrative purposes, we use the uniform and exponential distributions.
The efficacy as a function of sensor fidelity (x-axis) and the coupling (y-axis) is shown in the top panels of Fig. 4, with the light blue regions corresponding to those of high efficacy. The fully-connected graph is shown in the left-panel, while the Poisson and scale-free graphs are shown in the middle and right panels respectively. The width of the regions correlates with sensor fidelity, tapering as the amount of scrambled information increases. In the fully-connected case we see a strikingly different behavior in the low coupling regime, with the efficacy remaining high up until $\eta\approx 0.5$, as discussed in the last section. Both Poisson and scale-free graphs do not exhibit this high efficacy at low coupling, which we return to later. In all cases there is only a slight decrease in peak efficacy (dashed black curves in the top panels) with decreasing sensor fidelity. The high-efficacy region occurs for lower coupling strength in the scale-free case, but this is to be expected since $\sigma_{c}^{\textrm{scale-free}}<\sigma_{c}^{\textrm{Poisson}}$ as $\langle k^{2}\rangle^{\textrm{scale-free}}>\langle k^{2}\rangle^{\textrm{Poisson}}$.
Integrating over all coupling strengths, the viability (scaled to the actual viability), is shown in the middle and bottom panels of Fig. 4. In the middle panels a uniform distribution of couplings is used, while in the lower panels three exponential distributions are used. These differ in their means, chosen as $0.5\sigma_{c}^{A}$, $\sigma_{c}^{A}$, and $2\sigma^{A}_{c}$ with $\sigma_{c}^{A}\approx 0.5$. The solid lines represent simulations, and the dashed lines are linear fits. The figure indicates that the viability curves are not particularly sensitive to the choice of $\rho(\sigma)$.
The curves for the fully connected graph are best approximated by piece-wise linear fits indicating a transition in the viability per bit (VpB) from $1.5\pm 0.05$ at low fidelity to $0.25\pm 0.02$ at high fidelity. At poor sensor performance each bit of information is meaningful, significantly increasing system viability. This crossover indicates the presence of a semantic threshold. Once one crosses the threshold, the fidelity of the sensor becomes less important. This is analogous to prior work Sowinski et al. [2023], albeit with a steeper viability plateau.